OpenStudy (anonymous):

Simplify 17sqrt17-9sqrt17

OpenStudy (anonymous):

This is really like a kid's story problem. If you had 17 square roots of 17, and gave nine of them away, how many would you have left?

OpenStudy (anonymous):

\[17\sqrt{17}-9\sqrt{17}=8\sqrt{17}\]